OpenStudy (anonymous):

Translate the sentence into an equation using n as the unknown number. Then solve the equation for n. 5 increased by half a number is 11. a. 5 + 0.5n = 11 n = 12 c. 5(0.5n) = 11 n = 4.4 b. 5n + 0.5 = 11 n = 2.1 d. 5n +0.5 = 11 n = 12

OpenStudy (anonymous):

Just plug in the n. 5(0.5*n)=11 5(0.5*4.4)=11 5(2.2)=11 5*2.2=11